OpenStudy (anonymous):

if the temperature on the ground is 75°f,then the air temperature x miles high is given by T=75-29x.determine the altitudes at which the air temperature is less then 4°f. the altitudes at which the air temperature is less then 4°f are more than___mi.

OpenStudy (anonymous):

so here u need to find x given T<=4 hence 75-29x <= 4 or -29x <= 4-75 or -29x<=-71 or x>= 71/29 x>=2.448 appx

OpenStudy (anonymous):

2.4